A man is 46 years old and his grandson is 10 years old. In how many years will the godfather be thrice as old as his godson?
Answer by checkley77(12844)

46+X=3(10+X)
46+X=30+3X
X-3X=30-46
-2X=-16
X=-16/-2
X=8 YEARS GODFATHER WILL BE 3 TIMES THE GRANDSON'S AGE.
PROOF:
46+8=3(10+8)
54=3*18
54=54